Find an equation where ax^2+bx+c=0 with solutions of 2/5 and -5. Could you please answer within the hour? Thank you. A calculus teacher thought this would be fun homework worth a test grade. Please help!!!
Answer by richard1234(7193)

Just saying, this is nowhere near calculus level.

A fundamental rule of polynomials is that if r is a root of the polynomial, then x-r is a factor of the polynomial. The polynomial with roots 2/5 and -5 can be expressed as (x - 2/5)(x + 5) = x^2 + 5x - (2/5)x - 2 = x^2 - (23/5)x - 2. We may multiply this polynomial by some arbitrary non-zero constant C (preferably 5 to clear the fractional coefficient).